OpenStudy (czarluc):

What are the certain rules/steps in getting the domain of the operation of functions? namely sum,difference,product and quotient of functions?

OpenStudy (anonymous):

for the sum, difference and product... the domain will be the common domain, common to both functions. so if f + g, then the domain will be the common domain of both f and g; that is, the domain of f + g will be the same as all of the points that are common to the domain of f and the domain of g.

OpenStudy (anonymous):

for the quotient, f/g it's still the common domain with the added bit that you have to exclude whatever makes g (the denominator) equal 0. this is because division by 0 is not allowed.

OpenStudy (czarluc):

so just get the intersection of both functions for sum, diff and product?

OpenStudy (anonymous):

the intersection of the domains

OpenStudy (czarluc):

but what if the sum of the functions or diff or product. forms a rational function in the end? would I get the domain of that too?

OpenStudy (anonymous):

well, if the sum, diff or product forms a rational function, one of the original functions had to be a rational function. there would already be restriction on the original function so that there was no division by 0.

OpenStudy (czarluc):

f(x)=x^2-x-6 g(x)=x+2; find f(x)/g(x) =>(x+2)(x-3)/x+2 =>x-3

OpenStudy (czarluc):

what would the domain be?

OpenStudy (czarluc):

oh ok got it thanks... only confused in the domain for quotient

OpenStudy (anonymous):

so since f and g are both polynomials, their domains are all real numbers. but because you are dividing by g, you need to find if g can be 0. if it can, then that point in g's domain must be excluded from the domain of f/g.

OpenStudy (anonymous):

so what makes g equal 0?

OpenStudy (czarluc):

so the domain would be D={x/x not -2} ? even if the resulting function isn't a rational function?

OpenStudy (czarluc):

even if you canceled it in the solution?

OpenStudy (anonymous):

{\(x|x \in \mathbb{R},\,x \ne -2\)}

OpenStudy (anonymous):

yes, even if it cancels out. the fact is that you start with f/g and x = -2 must be excluded.

OpenStudy (anonymous):

what happens is you get a "hole" at the point x = -2.

OpenStudy (czarluc):

ohhhhh, thanks I'm gonna survive that test now haha! thank you for youe replies and patience..... :D

OpenStudy (anonymous):

you're welcome! thanks for sticking with it!

OpenStudy (czarluc):

*your